Circular No. 6741

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

PSR 0950+08 H. O. Vats, A. K. Singal, M. R. Deshpande, C. R. Shah, and S. Doshi,

Physical Research Laboratory, Ahmedabad; and K. N. Iyer and R. Oza, Saurashtra University, Rajkot, report a marked increase in the intensity level of the average-pulse amplitude of PSR 0950+08 during the first half of August in 103-MHz observations made with the 5000-m^{2} transit radio telescope at Rajkot. The average intensity, which peaked on Aug. 8, attained a level that was more than an order of magnitude higher than the normal average-pulse amplitude. At least 100 giant pulses were seen with an amplitude exceeding 100 times the average-pulse amplitude during a half-hour of transit time. Some of the strongest pulses reached a level of ∼ 500 times the average-pulse amplitude.

COMET C/1995 O1 (HALE-BOPP) D. H. Wooden, Ames Research Center, NASA; D. E. Harker and C. E.

Woodward, University of Wyoming; and H. A. Butner, University of Arizona, report 10-μm spectrophotometry (3′′ aperture; resolution ∼ 220) of comet C/1995 O1, obtained with the Wyoming Infrared Observatory 2.34m telescope (+ NASA/Ames HIFOGS) on Aug. 23.70 UT: “A weak silicate feature (9.0–11.6 μm) was detected with a peak flux density at 10.5 μm of 6.3 ± 1.7 × 10^{−18} W cm^{−2} μm^{−1}, with representative continuum flux densities at 8.3 and 11.8 μm of (2.5 ± 1.4) and (3.5 ± 1.6) × 10^{−18} W cm^{−2} μm^{−1}, respectively. Subsequent spectrophotometric measurements on Aug. 28.58 achieved the same spectral flux densities and shape within the photometric uncertainties. The weak silicate feature detected on Aug. 23 and 28 is 20 times less than that predicted using an r^{−4}Δ^{−1} extrapolation (Gehrz and Ney 1992, Icarus 100, 162) of the observed silicate-feature flux density (4.4 ± 0.1 × 10^{−16} W cm^{−2} μm^{−1} at 10.0 μm in a 3′′ aperture) on June 25.84, when the continuum was well fitted by a blackbody with a color temperature of 295 K with a continuum flux density at 8.6 μm of 1.8 ± 0.1 × 10^{−16} W cm^{−2} μm^{−1}.”

Total visual magnitude estimates: Aug. 8.81 UT, 3.4 (D. A. J. Seargent, The Entrance, N.S.W., naked eye); 13.79, 4.0 (Seargent); 20.34, 4.2 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars; 25.34, 4.2 (Aguiar); 31.34, 4.6 (R. Lourençon, Jundiai, Brazil, 7×50 binoculars); Sept. 3.33, 4.4 (W. C. de Souza, São Paulo, Brazil, 11×80 binoculars); 8.33, 4.6 (de Souza); 11.32, 4.6 (Aguiar).

1997 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green